OpenStudy (anonymous):

write an equation of the line with slope 4/5 and y intercept -2.

OpenStudy (anonymous):

y = 4/5 x - 2

OpenStudy (anonymous):

y=mx+b y=4/5x-2 do we stop here or do we have to say 5y -4x = -2/5

OpenStudy (anonymous):

I think stopping at slope-intercept form is enough.

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

You're welcome.